OpenStudy (anonymous):

Find the distance between the points given. (0, -6) and (9, 6) 15 √(26)

OpenStudy (anonymous):

15?

OpenStudy (danjs):

The distance formula

OpenStudy (danjs):

square root of [(x2-x1)^2 + (y1-y1)^2]

OpenStudy (danjs):

Just an application of the pythagorean theorem, you are measuring the length of the hypotenuse of a triangle, defined by those 2 points

OpenStudy (danjs):

and vertical and horizontal lines

OpenStudy (danjs):



OpenStudy (danjs):

\[distance = \sqrt{(y2-y1)^2 + (x2-x1)^2}\]

OpenStudy (danjs):

(x1,y1) = (0 , -6) (x2,y2) = (9, 6)

OpenStudy (danjs):

you good now @hhelpplzzzz ?

OpenStudy (danjs):

oh , not sure didnt calculate.. one sec

OpenStudy (solomonzelman):

sqrt(81+36) no?

OpenStudy (solomonzelman):

the other way, but makes no difference

OpenStudy (danjs):

i got 15

OpenStudy (solomonzelman):

oh lol its 12^2 inside, what a mistake

OpenStudy (danjs):

sqrt( 144 + 81)

OpenStudy (solomonzelman):

144+81 = 225, yes 15 is right

OpenStudy (danjs):

-6-(-6) = 6

OpenStudy (anonymous):

oh ok thnx

OpenStudy (danjs):

your welcome,